I’m working on quantitative numbers in chemistry. I’m confused on a problem.
I was given N=5. I was told to find from which outer orbital can you find the wave length of 4674
1. N=5, l=2
1
Expert's answer
2020-09-30T04:18:07-0400

c=lxv

l=c/v = 3x108/4674 = 64 184,8524